# Deconvoluting Thermomechanical Effects in X-ray Diffraction Data using Machine Learning

### v1

@@ 6 Discussion
Here we demonstrated a machine-learning-enabled approach to analyzing complex diffraction patterns from volumes containing distributions of thermal and mechanical strains. The work is an extension of a previous effort using a similar approach that was only able to extract information about distributions of temperature (thermal strain), with the presence of mechanical strain being a source of error. While the approach was applied to analyzing material being laser melted, it can also be applied to any rapid, complex testing scenario where distributions of thermal and mechanical strains are expected (e.g., dynamic loading). In addition, this type of approach provides a path forward for analyzing other types of lattice deformation that are composed of different types of Eigenstrains besides temperature (e.g, intrinsic piezoelectric or chemical).
@@
The ML approach here provides several primary benefits over more standard methods that analyze the shifts of diffraction peaks using analytical function fitting to probe temperature [12, 13, 14] or temperature and mechanical strain [41, 18]. First, the lack of peak fitting allows a wider range of peak shapes to be quantitatively analyzed. For example, ‘split’ diffraction peaks that are often encountered due to the wide differences, or distributions of temperatures, between inside and outside a melted region. Second, the thermomechanical modeling provides a means of decoupling thermal strains and volumetric elastic strains which will deform the lattice in the same fashion. Third, the method provides metrics about the distributions of strains present which can be of significant interest for microstructural evolution as many features develop as function of field gradients. Lastly, the GPR surrogate models provide a natural measure of uncertainty through output of the variance of the model predictions.
@@
As was seen in §4, the GPR models were most accurate at predicting the mean of the thermal and elastic strain distribution present in the reserved testing data. This is not surprising as the mean of the distribution also contributes to the peak of diffracted intensity. Importantly, the GPR models appear to be effective at fully deconvoluting thermal and elastic strains which, again, is a challenge due to the inability to directly differentiate between thermal and volumetric elastic strains in cubic materials; applying the GPR models to the in situ experimental data yielded evolution of thermal and elastic strain that were consistent with physics-based modeling.
@@
While still generally accurate, the predictions of maximum and minimums of the various strain distributions were subject to under or over predictions, usually at the extremes of the ranges of strain used for model training. The predictions of standard deviations of strain distributions were also generally underpredicted. Both of these observations are likely due to the fact that volumes of material with extreme values of strain generally do not contribute to the primary body of diffraction peaks. Further effort is required to determine if these issues could be addressed by increasing the amount of training data. In particular, increasing the number of microstructural configurations used for GPR model training may be beneficial (i.e., instantiating the same thermomechanical distributions with different grain orientation configurations).
@@
For this effort, we took the approach of performing high-fidelity simulations for both generating thermal (finite difference heat transfer and fluid flow) and elastic (finite element elasto-plasticity) strain distributions. The rationale of this approach was that by training with high-fidelity thermal and elastic strain fields that may be present, less training data would be necessary to analyze experimental data from similar conditions. This approach appears to be successful, providing results from experimental data in line with what is expected, but the downside is that the total modeling of the physical melting and cooling process is challenging. However, another approach may be to train the models with high volumes of low-fidelity simulation data. This would ease the accessibility and adoption of our proposed approach, but requires further research to evaluate feasibility. Similarly, an avenue that requires further investigation is the sensitivity of the GPR model predictions to material parameters used in the thermomechanical modeling for GPR training. While the GPR models are effectively connecting thermal and elastic strain distributions with diffracted intensity distributions, material parameters will alter the space of distributions for model training, which may alter final predictions. Again, we opted for use of accurate material parameters, specifically those generated with first-principles DFT. However, this process is intensive and our presented approach would be significantly eased if literature values with less provenance can achieve similar prediction results.
@@ 7 Summary
High-fidelity atomistic, heat-transfer and fluid flow, thermomechanical, and X-ray scattering simulations were brought together to train machine-learning (Gaussian Process Regression, GPR) surrogate models for analyzing diffraction data with complex thermomechanical distributions of deformation present. The trained GPR models were then transferred into the experimental domain to extract thermal and elastic strain distribution metrics from in situ diffraction data gathered during laser melting of IN625. In summary:
@@
    The physical modeling was used to generate over 2000 diffraction images with various thermal, elastic, and microstructural configuration for GPR model training.
@@
    The GPR models are capable of separating the effects of thermal and mechanical elastic strain distributions, including at high temperatures where the magnitude of thermal stains are significantly larger.
@@
    The GPR models were most accurate at predicting the mean of the thermal and elastic strain distributions from diffraction data, followed by the maximum and minimums of the distribution. In general, the GPR models tended to under predict the standard deviation (or spread) of the thermal and elastic strains present.
Acknowledgements
@@
All data used for this work is available upon reasonable request. The Python-based diffraction simulation and GPR training codes are also available upon request.

### v3

@@ 6 Discussion
Here, we demonstrated a machine-learning-enabled approach to analyzing complex diffraction patterns from volumes containing distributions of thermal and mechanical strains. Future application of this approach is to support the development of new classes of process diagnostics for (additive) manufacturing. Porosity from lack-of-fusion or keyholing and residual stress from unexpected process excursions are still challenges leading to failed builds. Surface optical measurements and thermography provide some insight into porosity formation, but in particular don’t provide insight into residual stress development. While not immediate, rapid advances in high-brightness laboratory X-ray source technology (e.g., current metal jet [33] and future thin-film diamond sources [34, 35]) could lead to in-chamber AM diagnostics. To characterize the thermomechanical state of the material in-chamber using these sources, non-standard analysis of diffraction data is needed. As opposed to trying to determine what the average material response is under relatively well-characterized thermomechanical loading, instead, a process diagnostic needs to quantify thermomechanical state using relatively well-understood material response (which is the approach taken here). While our approach was applied to analyzing material being laser melted, it can also be applied to any rapid, complex testing scenario where distributions of thermal and mechanical strains are expected (e.g., dynamic loading). In addition, our approach provides a path forward for analyzing other types of challenging material processing scenarios where measured lattice deformation is composed of different types of eigenstrains besides temperature (e.g., intrinsic piezoelectric or chemical).
@@
Using the ML surrogate model trained with thermomechanical simulations for analyzing the diffraction data provides several primary benefits over more standard methods that analyze the shifts of diffraction peaks using analytical function fitting to probe temperature [12, 13, 14] or temperature and mechanical strain [36, 18]. First, the lack of peak fitting allows a wider range of peak shapes to be quantitatively analyzed. For example, ‘split’ diffraction peaks are often encountered due to the wide differences, or distributions, of temperatures between the inside and outside of a melted region. Second, the thermomechanical modeling provides a means of decoupling thermal strains and volumetric elastic strains which will deform the lattice in the same fashion. Third, the method provides metrics about the distributions of strains present which can be of significant interest for microstructural evolution as many features develop as functions of field gradients. Lastly, the GPR surrogate models provide a natural measure of uncertainty through output of the variance of the model predictions.
@@
As was seen in §4, the GPR models were most accurate at predicting the mean of the thermal and elastic strain distributions present in the reserved testing data. This is not surprising as the mean of the distribution also contributes to the peak of diffracted intensity. Importantly, the GPR models appear to be effective at fully deconvoluting thermal and elastic strains which, again, is a challenge due to the inability to directly differentiate between thermal and volumetric elastic strains in cubic materials; applying the GPR models to the in situ experimental data yielded evolution of thermal and elastic strains that were consistent with physics-based modeling.
@@
While still generally accurate, the predictions of maximums and minimums of the various strain distributions were subject to under- or over-predictions, usually at the extremes of the ranges of strain used for model training. The predictions of standard deviations of strain distributions were also generally under-predicted. Both of these observations are likely due to the fact that volumes of material with extreme values of strain generally do not contribute to the primary body of diffraction peaks and these extreme thermomechanical states are also extremely transient and as such have a small presence in the simulated training data. Further effort is required to determine if extreme value accuracy could be improved by increasing the number of simulations and the amount of data used for surrogate model training. In particular, increasing the number of microstructural configurations used for GPR model training may be beneficial (i.e., instantiating the same thermomechanical distributions with different grain orientation configurations).
@@
For this effort, we took the approach of performing high-fidelity simulations for both generating thermal (finite difference heat transfer and fluid flow) and elastic (finite element elasto-plasticity) strain distributions. The rationale of this approach was that by training with high-fidelity thermal and elastic strain fields that may be present, less training data would be necessary to analyze experimental data from similar conditions. This approach appears to be successful, providing results from experimental data in line with what is expected, but the downside is that the complete modeling of the physical melting and cooling process is challenging. To lower the modeling burden, an open question worth investigating is what fidelity of thermomechanical simulations are necessary for sufficiently accurate thermomechanical quantification from experimental measurements? Another approach may be to train the models with lower computational burden semi-analytical models [37] or high volumes of low-fidelity simulation data. This would ease the accessibility and adoption of our proposed approach. Similarly, another avenue that requires further investigation is the sensitivity of the GPR model predictions to microstructural and material parameters used in the thermomechanical modeling for GPR training. While the GPR models are effectively connecting thermal and elastic strain distributions with diffracted intensity distributions, microstructural and material parameters will alter the space of distributions for model training, which may alter final predictions. In this work, the morphologies and size distributions of grains is not a perfect match to experiment, but the GPR model still produces rational results indicating that model accuracy may be able to be relaxed. Furthermore, here we opted for use of accurate material parameters, specifically those generated with first-principles DFT. However, this process is intensive and our presented approach would be significantly eased if literature values with less provenance or similar alloy compositions (e.g., using pure Ni properties for a superalloy) could achieve similar prediction results.
@@ 7 Summary
High-fidelity atomistic, heat transfer and fluid flow, thermomechanical, and X-ray scattering simulations were brought together to train machine-learning (Gaussian Process Regression, GPR) surrogate models for analyzing diffraction data with complex thermomechanical distributions of deformation present. The trained GPR models were then transferred into the experimental domain to extract thermal and elastic strain distribution metrics from in situ diffraction data gathered during laser melting of IN625. In summary:
@@
    The physical modeling was used to generate over 2000 diffraction images with various thermal, elastic, and microstructural configurations for GPR model training.
@@
    The GPR models are capable of separating the effects of thermal and mechanical elastic strain distributions, including at high temperatures where the magnitude of thermal strains are significantly larger.
@@
    The GPR models were most accurate at predicting the mean of the thermal and elastic strain distributions from diffraction data, followed by the maximums and minimums of the distribution. In general, the GPR models tended to under-predict the standard deviation (or spread) of the thermal and elastic strains present.
Acknowledgments
@@
All data used for this work are available upon reasonable request. The Python-based diffraction simulation and GPR training codes are also available upon request.